The perimeter of a rectangle is 32 ft. The length is 2 ft. longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width.
What is the length ? ft.
What is the width ? ft.
Could someone please help me I would appreciate it. Thanks

Answer by scott8148(6628)

"The perimeter of a rectangle is 32 ft" ___ 2x + 2y = 32

"The length is 2 ft. longer than the width" ___ x = y + 2 ___ 2x - 2y = 4

adding equations ___ 4x = 36 ___ x = 9

substituting ___ (9) = y + 2 ___ 7 = y